# Learning Structured Representations by Embedding Class Hierarchy with Fast Optimal Transport

### v1

@@ G.1 Effect of Tree Depth
G.2 Effect of Tree Weights
@@
In this section, we demonstrate that all CPCC methods, including OT-CPCC, are influenced by input tree weights. In Fig.11, we begin with the original unweighted CIFAR100 tree (where all edges have a default weight of 1) and modify the edge weights of a subtree representing the coarse label aquatic_mammals, which includes five fine labels: beaver, dolphin, otter, seal, and whale. We visualize the learned representation through a distance matrix, as described in Sec.G.1. From left to right in Fig.11, the edge weights of this subtree are set to 1, 2, and 4, respectively. Consequently, we observe a color shift in the top-left coarse block of the matrix, with the block gradually darkening as the subtree weights increase. This demonstrates the flexibility of OT-CPCC methods, which can successfully embed weighted trees, and highlights why CPCC is a sensitive and effective metric for evaluating the interpretability and quality of the hierarchical information embedded in the learned representation, as discussed in Sec.4.4.
G.3 Effect of Flow Weight
@@
G.4 Comparison to Other Hierarchical Baselines
@@
In Fig.12, we compared OT-CPCC methods with other hierarchical objectives described in Sec.F, and we report the Fine/Coarse Metric Avg., which averages all source/target classification and retrieval metrics. Based on CPCC, fine, and coarse metrics, we observe that OT-CPCC methods behave similarly to $\ell_{2}$-CPCC, exhibiting higher CPCC values and stronger performance on coarse metrics, but lower fine metrics (though OT methods still outperform $\ell_{2}$ on fine metrics, as shown in Tab.5). In contrast, all other hierarchical baseline objectives perform more like flat cross-entropy methods, with the opposite trend—lower CPCC and coarse metrics but relatively higher fine metrics. Additionally, we excluded Rank loss from the correlation plot on the right side of Fig.12, as Tab.18 suggests that Rank loss is not suitable for any of the downstream tasks used in our evaluation, making it an outlier compared to the other objectives.
@@
Appendix H Implementation Details for Synthetic Datasets Experiments

### v2

@@ H.1 Effect of Tree Depth
Table 13: EMD and FastFT performance for tree with different depths on the full CIFAR100 dataset. The batch size is set to $128$.
Objective Setup CPCC FineAcc CoarseAcc
EMD depth = 3 88.80 71.91 82.21
depth = 4 85.09 72.59 82.93
depth = 5 88.64 72.20 82.50
FastFT depth = 3 90.75 72.71 83.31
depth = 4 87.86 72.45 83.05
depth = 5 90.27 72.56 83.04
We also include metrics for the preciseness of hierarchical information and hierarchical classification in Tab.13. As observed in Fig.10, the tree information is well-learned, leading us to expect high CPCC scores across all settings in Tab.13. For the hierarchical classification metrics, within the same OT method, the difference between FineAcc and CoarseAcc is small. It is worth noting that, because we train on the full CIFAR100 dataset for visualization purposes, the source-target hierarchical split shown in Fig.4 is not applied here. Both FineAcc and CoarseAcc are reported on the original CIFAR100 test set. This suggests that even when the tree used is not identical to the label hierarchy associated with the dataset, an approximate tree with a similar structure is sufficient for classification tasks.
@@
H.2 Effect of Tree Weights
@@
In this section, we demonstrate that all CPCC methods, including OT-CPCC, are influenced by input tree weights. In Fig.11, we begin with the original unweighted CIFAR100 tree (where all edges have a default weight of 1) and modify the edge weights of a subtree representing the coarse label aquatic_mammals, which includes five fine labels: beaver, dolphin, otter, seal, and whale. We visualize the learned representation through a distance matrix, as described in Sec.H.1. From left to right in Fig.11, the edge weights of this subtree are set to 1, 2, and 4, respectively. Consequently, we observe a color shift in the top-left coarse block of the matrix, with the block gradually darkening as the subtree weights increase. This demonstrates the flexibility of OT-CPCC methods, which can successfully embed weighted trees, and highlights why CPCC is a sensitive and effective metric for evaluating the interpretability and quality of the hierarchical information embedded in the learned representation, as discussed in Sec.4.4.
Table 14: EMD and FastFT performance for tree with different edge weights within aquatic_mammals classes on the full CIFAR100 dataset. The batch size is set to $1024$.
Objective Setup CPCC FineAcc CoarseAcc
EMD weight 1x 89.13 75.76 85.50
weight 2x 88.85 75.57 85.22
weight 4x 79.86 75.56 85.35
FastFT weight 1x 92.12 75.43 85.43
weight 2x 92.57 75.38 85.10
weight 4x 84.23 75.57 84.98
We use the same experimental settings as those described in Sec.H.1. While the classification metrics remain comparable, the most significant trend emerges in the CPCC metric. As shown in Fig.11, we observe the expected pattern in the top-left block, which changes as the edge weights are multiplied. However, as the injected hierarchy deviates further from the given label hierarchy, it becomes increasingly unnatural. The increased weights make the tree embedding more challenging, resulting in a drop in CPCC values for larger weights.
H.3 Effect of Flow Weight
@@
Appendix I Analysis for the Time Complexity of OT approximation methods
In this section, we provide the details and references for the reported time complexities in Tab.1.
In the following time complexity analysis, denote $b$ as batch size. We assume each class-conditioned cluster has the same size $n$, which implies $b=nk$, where $k$ is the number of classes.
For $\ell_{2}$, given one batch of batch size $b$, computing Euclidean centroids for all classes requires scanning through all $d$-dimensional data points once, which requires $O(bd)=O(nkd)$. We have at most $k$ classes in one batch, so computing the pairwise distances between class centroids will cost $O(k^{2}d)$. The total cost for $\ell_{2}$ is $O((nk+k^{2})d)$.
[…]
For EMD, the complexity is typically $O(n^{3}\log n)$ in the past literature (Cuturi, 2013), which primarily focuses on the influence of the number of samples $n$. However, in our setting, because $d>>n$, we aim to emphasize the effect of $d$. In the time complexity expression for EMD, the $O(n^{3}\log n)$ term arises from solving the linear programming problem, while the $O(nd)$ term accounts for computing the objective function ${\bm{P}},{\bm{D}}\right>$. Due to the sparse nature of the linear programming solution, there are only $O(n)$ non-zero values in ${\bm{P}}$. Consequently, we compute $O(n)$ $d$-dimensional pairwise distances in ${\bm{D}}$, making the dot product computation require $O(nd)$ in total. Therefore, the overall complexity of EMD is $O(n(d+n^{2}\log n))$.
To get the approximated flow matrix, Sinkhorn (Cuturi, 2013) mainly depends on the iterative matrix scaling algorithm. For each $I$ iteration, we normalize the row and column of a matrix of size $O(n^{2})$, so the complexity for running this iterative algorithm is $O(n^{2}I)$. Unlike in EMD, the approximated flow matrix $P$ derived from Sinkhorn is not sparse. Therefore, the dot product of the objective function will now take $O(n^{2}d)$ in total. The total complexity of Sinkhorn is $O(n^{2}(d+I))$.
To obtain the approximated flow matrix, Sinkhorn (Cuturi, 2013) relies on the iterative matrix scaling algorithm. During each of the $I$ iterations, we normalize the rows and columns of a matrix of size $O(n^{2})$, resulting in a complexity of $O(n^{2}I)$ for the iterative algorithm. Unlike EMD, the approximated flow matrix ${\bm{P}}$ derived from Sinkhorn is not sparse. Consequently, computing the dot product for the objective function requires $O(n^{2}d)$ in total. Therefore, the overall complexity of Sinkhorn is $O(n^{2}(d+I))$.
For FT (Backurs et al., 2020), the most computationally expensive step is the tree construction using the QuadTree algorithm. Backurs et al. (2020) demonstrate that this tree can be built in $O(nd\log(d\Phi))$. Once the tree is constructed, as shown in Lemma 3.1 of Backurs et al. (2020), FT requires $O(n(d+\log(d\Phi)))$ to run the bottom-up algorithm and compute the objective function, which is less expensive than the tree construction. Thus, the total complexity of FT is $O(nd\log(d\Phi))$.
For Fast FlowTree, please refer to the proof of Thm. 3.2.
For TWD, we follow Algorithm 1 in Yamada et al. (2022). The first step is to construct a tree. To save computation, instead of using QuadTree, we leverage the same technique as in FastFT to use our augmented tree, making this step $O(1)$. For each of the $I$ iterations in supervised learning, we need to compute all entries in the pairwise distance matrix ${\bm{D}}$, which requires $O(n^{2}d)$. Thus, even when the tree structure is known, the total complexity of TWD is $O(n^{2}dI)$.
Finally, we want to highlight that Fig.3 is evaluated with $k=2$, where we vary the number of samples in one source and one target distribution. In this scenario, although the theoretical computational complexities for $\ell_{2}$ and Fast FlowTree are both $O(nd)$, we observe that $\ell_{2}$ scales better in practice. This is due to the effect of standard vectorization optimizations in common matrix operation packages, which leverage parallelism along the dimension of $n$.
I.1 Comparison to Other Hierarchical Baselines
@@
In Fig.12, we compared OT-CPCC methods with other hierarchical objectives described in Sec.G, and we report the Fine/Coarse Metric Avg., which averages all source/target classification and retrieval metrics. Based on CPCC, fine, and coarse metrics, we observe that OT-CPCC methods behave similarly to $\ell_{2}$-CPCC, exhibiting higher CPCC values and stronger performance on coarse metrics, but lower fine metrics (though OT methods still outperform $\ell_{2}$ on fine metrics, as shown in Tab.5). In contrast, all other hierarchical baseline objectives perform more like flat cross-entropy methods, with the opposite trend—lower CPCC and coarse metrics but relatively higher fine metrics. Additionally, we excluded Rank loss from the correlation plot on the right side of Fig.12, as Tab.20 suggests that Rank loss is not suitable for any of the downstream tasks used in our evaluation, making it an outlier compared to the other objectives.
@@
Appendix J Implementation Details for Synthetic Datasets Experiments
@@
Appendix K Comparison of Runtime of OT-CPCC on Real-world Datasets
Table 21: Comparison of the wall clock runtime of OT-methods on real datasets for training one epoch. The results are averaged for three random seeds.
Dataset Objective Per Epoch (s) Dataset Objective Per Epoch (s)
Flat 85.88 (3.69) Flat 147.67 (4.54)
$\ell_{2}$ 83.78 (9.80) $\ell_{2}$ 149.39 (1.24)
FastFT 85.21 (4.26) FastFT 132.01 (3.24)
EMD 87.70 (2.43) EMD 234.68 (4.40)
Sinkhorn 85.65 (4.15) Sinkhorn 165.69 (4.60)
CIFAR10 (b=128) SWD 92.90 (4.65) INAT (b=1024) SWD 152.52 (8.90)
We include the wall clock run times for training different CPCC methods over one epoch in Tab.21, using two distinct settings: CIFAR10 with a batch size of 128 and model training from scratch, and INAT with a batch size of 1024 and model fine-tuning from an ImageNet checkpoint. Despite the differences in settings, we observe that FastFT consistently runs significantly faster than exact EMD computation and outperforms the other two approximation methods, SWD and Sinkhorn. This aligns with our theoretical analysis in Tab.1 and observations on synthetic data in Fig. 3. When comparing FastFT to $\ell_{2}$, we find that FastFT is not much slower, and in the case of INAT, it is even faster than $\ell_{2}$. Thus, we can confidently conclude that our proposed method, especially FastFT, is scalable.
Table 22: Comparison of the wall clock runtime of Bures-CPCC vs. OT-CPCC on real datasets for training one batch. The results are averaged for three random seeds.
Dataset Objective Per Batch (s)
$\ell_{2}$ 0.35 (0.18)
CIFAR100 EMD 1.94 (0.12)
Bures 33.77 (4.73)
Similar to $\ell_{2}$, one could argue for using another closed-form Wasserstein expression, the Bures-Wasserstein distance (as defined in Eq.7), in CPCC optimization. However, Bures depends on the covariance term of the source and target distributions, which requires at least $O(nd^{2})$ computation. As shown in Tab.1, in $\ell_{2}$, calculating the mean of a class-conditioned cluster only requires $O(nd)$, and all other OT methods have a linear $d$ term in their time complexity. In CIFAR100, even with a large batch size of 1024, $n$ is approximately $1024/100\approx 10$, assuming all classes are of equal size. Using ResNet18 as the backbone, $d=512$, which is typically much larger than $n$. Therefore, the $d^{2}$ factor from computing the covariance matrix in Bures can make it prohibitively expensive in our case. To verify this, we computed the running time for CIFAR100 (batch size = 128) for different methods in Tab.22, considering one batch update. The results are averaged over three random seeds.
As observed in Tab.22, Bures-Wasserstein is significantly slower than both $\ell_{2}$ and EMD due to its quadratic dependence on $d$. Therefore, for scalability concerns, covariance-based Wasserstein metrics are not well-suited to our problem setup.